# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v1

@@ Appendix B Applied Field and Internal Field
The power loss per cycle $Q_{\text{tot}}$ given by Eq. (4) can equivalently be computed in terms of $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\boldsymbol{m}/2$. Indeed, as also shown in [38],
@@
That is, the areas enclosed by the loops represented in Figs. 5 and 7 are identical.
Appendix C Hysteresis Model Jacobian
@@
Its expression can be computed independently for each cell. Below, we give its expression for the simple case of a S cell and then generalize for a CS cell. We drop the index $k$ for conciseness.
@@ C.1 S cell
with $\mathbf{I}$ the second-order identity tensor and $\mathbf{K}$ defined by the following dyadic product:
$\mathbf{K}=(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}).$ (50)
@@ C.2 CS cell
By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (35), (33), and (34), which yields the following reversible field
@@
with
$\mathbf{L}=(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}).$ (56)

### v2

@@ Appendix B Applied Field and Internal Field
The energy loss per cycle $Q_{\text{tot}}$ given by Eq. (4) can be equivalently computed in terms of $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\boldsymbol{m}/2$. Indeed, as also shown in [36],
@@
In particular, this justifies that the areas enclosed by the loops represented in Figs. 4 and 6 are identical.
Appendix C ROHM model Jacobian
@@
Its expression can be computed independently for each cell. Below, we give its expression for the simple case of a S cell and we then generalize for a CS cell. We drop the index $k$ for conciseness.
@@ C.1 S cell
with $\mathbf{I}$ the second-order identity tensor and $\mathbf{K}$ the dyadic product $\mathbf{K}=(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})$.
@@ C.2 CS cell
By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (38), (36), and (37), which yield the following reversible field
@@
with $\mathbf{L}=(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})$.